# Software Engineering and Foundation Models: Insights from Industry Blogs Using a Jury of Foundation Models

### v1

@@ Abstract
Foundation models (FMs) such as large language models (LLMs) have significantly impacted many fields, including software engineering (SE). The interaction between SE and FMs has led to the integration of FMs into SE practices (FM4SE) and the application of SE methodologies to FMs (SE4FM). While several literature surveys exist on academic contributions to these trends, we are the first to provide a practitioner’s view. We analyze 155 FM4SE and 997 SE4FM blog posts from leading technology companies, leveraging an FM-powered surveying approach to systematically label and summarize the discussed activities and tasks. We observed that while code generation is the most prominent FM4SE task, FMs are leveraged for many other SE activities such as code understanding, summarization, and API recommendation. The majority of blog posts on SE4FM are about model deployment & operation, and system architecture & orchestration. Although the emphasis is on cloud deployments, there is a growing interest in compressing FMs and deploying them on smaller devices such as edge or mobile devices. We outline eight future research directions inspired by our gained insights, aiming to bridge the gap between academic findings and real-world applications. Our study not only enriches the body of knowledge on practical applications of FM4SE and SE4FM but also demonstrates the utility of FMs as a powerful and efficient approach in conducting literature surveys within technical and grey literature domains. Our dataset, results, code and used prompts can be found in our online replication package at https://github.com/SAILResearch/fmse-blogs.
@@ IV-A Identifying blogs related to SE and FM from technology companies
TABLE II: Cohen’s Kappa between LLMs and human labels on the golden set for SE-FM area, SE4FM activity, and FM4SE activity classification tasks.
@@
FM4SE
Activities
Gemini-1.5-Flash-002 0.65 0.81 0.66
GPT-4o-mini-2024-07-18 0.70 0.81 0.74
Qwen2-72B-Instruct 0.85 0.76 0.79
LLM Jury 0.95 0.91 0.87
@@ IV-B Identifying the FM4SE and SE4FM area, activities and discussed tasks in the blog posts
We first label blog posts as FM4SE or SE4FM-related. Second, we label the FM4SE or SE4FM activities discussed in the blog posts. Finally, we summarize the activity-specific tasks that are discussed, to facilitate our manual review of the posts. Tables III and IV show the identified activities and tasks.
@@
Step 1 – Label blog posts as FM4SE or SE4FM with FM/LLM Jury. We used the FM/LLM Jury framework (Section III) to classify blog posts as SE4FM, FM4SE, or unrelated. We created a golden dataset of 100 randomly selected blog posts, manually labelled by the first author, to evaluate the FM/LLM Jury for this classification. Table II, shows that the FM/LLM Jury achieved an excellent agreement ($\kappa=0.95$) with human labels on the test set. We then applied the FM/LLM Jury to the entire dataset of 4,463 blogs, classifying 3,126 as unrelated, 156 as FM4SE and 1,122 as SE4FM.
@@
Step 2 – Label the FM4SE activities in the blog posts. We used the FM4SE activity labels from prior work [17]. A golden set of 30 FM4SE blogs was randomly sampled and manually labelled by the first author. Following our iterative prompt construction process (Section III), we reached an excellent agreement ($\kappa=0.91$) between the FM/LLM Jury and the human labels (see Table II). We used the best-performing prompt to label all FM4SE activities in the 156 FM4SE blog posts. One blog post was labelled as ‘Other’ as the FM/LLM Jury could not identify any specific FM4SE activities, ending up with 155 FM4SE blog posts in total.
@@
Step 3 – Label the SE4FM activities in the blog posts. Following a similar process, we labelled the SE4FM activities in the 1,122 SE4FM blogs. We created a list of SE4FM activity labels based on prior work [1, 28, 15]. We randomly sampled a golden set of 30 SE4FM blog posts and manually labelled the SE4FM activities. Following our prompt construction process, the FM/LLM Jury demonstrated strong agreement ($\kappa=0.87$) with the human labels. We used the best-performing prompt to label the SE4FM activities in all 1,122 SE4FM blog posts. 125 blog posts were labelled as ‘Other’ as the FM/LLM Jury could not identify any SE4FM activities, ending up with 997 FM4SE blog posts in total.
@@
Step 4 – Identify the activity-specific tasks in the blog posts. To facilitate our manual review of the blog posts, we prompted the FMs to take an additional step to identify the activity-specific tasks in the blog posts. For FM4SE blogs, we prompted the FMs to identify relevant tasks from a curated set of tasks [17, 12, 24]. Similarly, for SE4FM blogs, the FMs identified SE4FM-related tasks from a curated set of tasks [1, 28, 15]. If no predefined tasks were relevant, we asked the FMs to generate new tags based on the blog content. In cases where the FMs proposed new task tags, we employed a human-in-the-loop process to determine whether to reclassify or ignore these tags. However, no human intervention was required for FM4SE posts, and eight new SE4FM task tags identified by the FM model Gemini-1.5-Flash were resolved by the FM/LLM Jury.
@@ V RQ1: Which FM4SE activities are discussed in industry blog posts?
Software development 100 Code generation 11
@@ V-C Software maintenance
V-D Software management

### v2

@@ Abstract
Foundation models (FMs) such as large language models (LLMs) have significantly impacted many fields, including software engineering (SE). The interaction between SE and FMs has led to the integration of FMs into SE practices (FM4SE) and the application of SE methodologies to FMs (SE4FM). While several literature surveys exist on academic contributions to these trends, we are the first to provide a practitioner’s view. We analyze 155 FM4SE and 997 SE4FM blog posts from leading technology companies, leveraging an FM-powered surveying approach to systematically label and summarize the discussed activities and tasks. We observed that while code generation is the most prominent FM4SE task, FMs are leveraged for many other SE activities such as code understanding, summarization, and API recommendation. The majority of blog posts on SE4FM are about model deployment & operation, and system architecture & orchestration. Although the emphasis is on cloud deployments, there is a growing interest in compressing FMs and deploying them on smaller devices such as edge or mobile devices. We outline eight future research directions inspired by our gained insights, aiming to bridge the gap between academic findings and real-world applications. Our study not only enriches the body of knowledge on practical applications of FM4SE and SE4FM but also demonstrates the utility of FMs as a powerful and efficient approach in conducting literature surveys within technical and grey literature domains. Our dataset, results, code and used prompts can be found in our online replication package at https://zenodo.org/records/14563992.
@@ IV-A Identifying blogs related to SE and FM from technology companies
TABLE II: Cohen’s Kappa between LLMs and human labels on the golden set for SE-FM area, FM4SE activity, and SE4FM activity classification tasks.
@@
Gemini-1.5-Flash-002 0.64 0.81 0.69
GPT-4o-mini-2024-07-18 0.70 0.81 0.77
Qwen2-72B-Instruct 0.85 0.76 0.73
FM/LLM Jury 0.92 0.91 0.86
@@ IV-B Identifying the FM4SE and SE4FM area, activities and discussed tasks in the blog posts
We first label blog posts as FM4SE or SE4FM-related. Second, we label the FM4SE or SE4FM activities discussed in the blog posts. Finally, we summarize the activity-specific tasks that are discussed, to facilitate our manual review of the posts. Tables III and IV show the identified activities and tasks. The first author, a PhD graduate with 5 years of SE research experience, conducted all labelling and validated through discussions with co-authors.
@@
Step 1 – Label blog posts as FM4SE or SE4FM with FM/LLM Jury. We used the FM/LLM Jury framework (Section III) to classify blog posts as SE4FM, FM4SE, or unrelated. A golden dataset of 100 blog posts was randomly selected to evaluate this classification. Table II shows that the framework achieved an excellent agreement ($\kappa=0.92$) with human labels on the test set. We then applied the framework to the entire dataset of 4,463 blog posts, classifying 3,126 as unrelated, 156 as FM4SE and 1,122 as SE4FM.
@@
Step 2 – Label the FM4SE activities in the blog posts. We used the FM4SE activity labels from prior work [17] and a golden set of 30 FM4SE posts was randomly sampled for evaluation. Following our iterative prompt construction process (Section III), we reached an excellent agreement ($\kappa=0.91$) between the FM/LLM Jury and the human labels (see Table II). We applied the Jury to label all FM4SE activities in the 156 FM4SE blog posts. One blog post was labelled as ‘Other’ as the FM/LLM Jury could not identify any specific FM4SE activities, ending up with 155 FM4SE blog posts in total.
@@
Step 3 – Label the SE4FM activities in the blog posts. Following a similar process, we labelled the SE4FM activities in the 1,122 SE4FM blogs. We created a list of SE4FM activity labels based on prior work [1, 28, 15]. We randomly sampled a golden set of 30 SE4FM blog posts and manually labelled the SE4FM activities. Following our prompt construction process, the FM/LLM Jury demonstrated strong agreement ($\kappa=0.86$) with the human labels. We used the best-performing prompt to label the SE4FM activities in all 1,122 SE4FM blog posts. 125 blog posts were labelled as ‘Other’ as the FM/LLM Jury could not identify any SE4FM activities, ending up with 997 FM4SE blog posts in total.
@@
Step 4 – Identify the activity-specific tasks in the blog posts. To facilitate our manual review of the blog posts, we prompted the FMs to take an additional step to identify the activity-specific tasks in the posts. For FM4SE blogs, we prompted the FMs to identify relevant tasks from a curated task set [17, 12, 24]. Similarly, for SE4FM blogs, the FMs identified SE4FM-related tasks from a curated task set [1, 28, 15]. If no predefined tasks were relevant, we asked the FMs to generate new tags based on the blog content. In cases where the FMs proposed new task tags, we employed a human-in-the-loop process to determine whether to reclassify or ignore these tags. However, no human intervention was required for FM4SE posts, and eight new SE4FM task tags identified by the FM Gemini-1.5-Flash were resolved by the FM/LLM Jury.
@@ V RQ1: Which FM4SE activities are discussed in industry blog posts?
Software development 101 Code generation 11
@@ V-C Software maintenance
V-D Software management, requirement engineering, and software design